What is the slope of this line: 5x - 8y = 9?
Answer by checkley79(3341)

5x - 8y = 9
-8Y=-5X+9
Y=-5X/-8+9/-8
Y=5X/8-9/8
[Y=mX+b] IS THE LINEAR EQUATION WHERE m=THE SLOPE.
THEREFORE THIS SLOPE=5/8 ANS.